There are four consumers willing to pay the following amounts for haircuts:

A: $7 B: $2 C: $8 and D: $5

There are four haircutting businesses with the following costs;

Firm 1: $3 Firm 2: $6 Firm 3: $4 and Firm 4: $2

Each firm has the capacity to produce only one haircut. For efficiency how many haircuts can be 

given? Also find out how large will be the total surplus.


1
Expert's answer
2021-04-01T08:34:30-0400

For efficiency 4 haircuts can be given, because every consumer has high enough willingness to pay to cover the producers costs.

The total surplus is:

TS = (7 + 2 + 8 + 5) - (3 + 6 + 4 + 2) = 22 - 15 = $7.